In a student laboratory exercise, the wavelength of a 48,300 Hz ultrasound wave is measured to be .716cm. What is the site temprature in (k)??


1
Expert's answer
2021-10-05T10:07:24-0400

Find the speed:


"c=\lambda f=0.00716·48300=345.8\text{ m/s}."

The speed of sound depends on temperature in the following way:


"c=20.05\sqrt{t+273.15},\\\space\\
t=\bigg(\frac c{20.05}\bigg)^2-273.15=24°\text{C}."